### v1

# Improving Vision-Language-Action Models via Chain-of-Affordance

Abstract
Robot foundation models, particularly Vision-Language-Action (VLA) models, have garnered significant attention for their ability to enhance robot policy learning, greatly improving robot’s generalization and robustness. OpenAI’s recent model, O1, showcased impressive capabilities in solving complex problems by utilizing extensive reasoning chains. This prompts an important question: can robot models achieve better performance in multi-task, complex environments by reviewing prior observations and then providing task-specific reasoning to guide action prediction?
In this paper, we introduce Chain-of-Affordance (CoA), a novel approach to scaling robot models by incorporating reasoning in the format of sequential robot affordances to facilitate task completion. Specifically, we prompt the model to consider the following four types of affordances before taking action: (1) object affordance — what object to manipulate and where it is; (2) grasp affordance — the specific object part to grasp; (3) spatial affordance — the optimal space to place the object; and (4) movement affordance — the collision-free path for movement. By integrating this knowledge into the policy model, the robot gains essential context, allowing it to act with increased precision and robustness during inference. Our experiments demonstrate that CoA achieves superior performance than state-of-the-art robot foundation models, such as OpenVLA and Octo. Additionally, CoA shows strong generalization to unseen object poses, identifies free space, and avoids obstacles in novel environments.
1 Introduction
Recent advancements in Vision-Language-Action (VLA) models have shown that training with internet-scale data can empower end-to-end policy learning models to outperform non-VLA models. However, current approaches often rely heavily on high-level planning or task decomposition by off-the-shelf LLMs or VLMs, limiting models from developing implicit reasoning on their own. OpenAI’s recent O1 model has demonstrated that LLMs can improve performance on complex tasks through extensive reasoning chains. It is expected that this reasoning capability if adapted to robotic models, could enhance their action robustness and generalizability. Yet, the exploration of self-driven reasoning in robotics remains under-explored, highlighting an important frontier for future research.
In this work, we propose Chain-of-Affordance, namely CoA, a novel perspective on generalizing model reasoning at test-time, and leverage such generated reasoning to facilitate the policy learning process. Our model builds upon the DiffusionVLA [43], a state-of-the-art VLM model combining autoregression and diffusion objectives. Our method leverages visual affordance in robot learning, conceptualizing various actions and interactions with objects or the environment that a robot can perform based on visual contexts. We consider four types of affordance that are critical for robots to understand their observational surroundings to interact effectively with objects and achieve tasks in dynamic:
  • Object affordance. When a user provides vague instructions, the robot should be capable of identifying the target object for interaction and its location within the view.
  • Grasp affordance. It involves the robot assessing the object’s most appropriate points or surfaces to enable secure and stable handling.
  • Spatial affordance. The robot needs to identify a set of coordinates that satisfy relationships described in language, such as free space for placement.
  • Movement affordance. Identifying a trajectory for the robot to move without collision is crucial in the real world to avoid catastrophic damage.
These four affordances form a sequential chain, requiring the robot to possess prior knowledge at each step to advance to the next. Specifically, at test time, the robot must first understand ”what to manipulate and where the object is located.” Next, it determines ”how to grasp the object” and finally, ”where to place it,” all while following a trajectory that ensures safe task completion. During inference, the affordance chain is progressively generated as the action state evolves, avoiding unnecessary computational costs associated with outputting extensive language. We introduce two formats for chain-of-affordance reasoning: text-based and image-based chain-of-affordance prompting. Our model’s architecture natively supports text prompting integration, but image prompting requires adaptation. To address this, we developed an image affordance injection module that seamlessly integrates visual affordances into the policy network. Once generated by the model, this affordance knowledge is reused in policy learning, enabling the model to produce robust and generalizable actions.
We conduct extensive experiments on both simulated benchmarks and real-world tasks. Specifically, on the LIBERO [27] benchmark, our proposed CoA outperforms several other state-of-the-art approaches, including the Diffusion Policy [9], Octo [33], and OpenVLA [21]. Due to the substantial sim-to-real gap, many existing studies also evaluate real robots. We set up seven real-world robot tasks, including long-horizon and challenging tasks such as serving tea, cleaning garbage, and wiping tables. We perform multi-task learning on these real-world tasks and demonstrate that the VLA model, enhanced by chain-of-affordance reasoning, successfully handles most of these complex tasks with a high average success rate. Furthermore, CoA exhibits strong generalization capabilities to handle unseen object poses, adjust motion when obstacles appear, and identify free space for placement.
2 Related Works
Affordance in robotics. In robot learning, the concept of affordance is interpreted in various ways. Typically, affordance is defined as the functions of an object, encompassing what the object is, how it can be manipulated, and its spatial relationship to the target. This concept extends beyond visual properties, linking observations directly to actions. The efficacy of affordance prediction has been shown by many learning-based manipulation methods for 6-DoF grasping [12, 11, 39] and stable object placement. It can also represented in many ways [31, 30, 49, 46, 17, 25, 5, 28, 16] such as part segmentation, dense image feature descriptors and keypoints. Some methods leverage human videos to obtain affordances [2], while others use vision-language models (VLMs) to predict points representing spatial affordances [8, 47]. RT-Affordance [32] employs more descriptive affordance representations, and TraceVLA [1] incorporates visual traces as additional input to enhance VLA. Our approach introduces multiple types of affordances and formulates them as chain-of-thought reasoning to further improve VLA model performance.
Reasoning for language and control. Prompting large language models (LLMs) with ”think step-by-step” [42] has significantly advanced their ability to solve complex tasks. Numerous approaches [3, 50] have since been developed to encourage deeper reasoning in LLMs, such as Tree-of-Thought [45] and Chain-of-Code [23], establishing this as a standard practice in language modeling. Recent research has leveraged LLMs and vision-language models (VLMs) [15, 13, 25, 10, 16, 14] as high-level planners in robotics, often using fine-tuned, open-source models or closed-source LLMs alongside policy networks for low-level task execution. These studies illustrate that detailed reasoning can enhance low-level control. ECoT [48] introduces a reasoning strategy for VLMs that includes task decomposition, subtask descriptions, fine-grained movement instructions, gripper positioning, and object tracking on the table. However, the extensive output increases computational costs. Our work proposes a more concise and efficient reasoning approach, improving low-level control while reducing computational demands.
3 Preliminary on Vision-Language-Action Models
Our work builds upon Vision-Language-Action (VLAs) as the backbone for our chain-of-affordance policies. VLAs employ a straightforward policy-learning approach: beginning with a pre-trained vision-language model, they fine-tune it to predict the next robot action $a$ based on the current image observation $I$, task instruction $T$, and reasoning $r$. There are two types of VLAs: autoregressive VLAs [7, 34, 1, 48, 21] and diffusion-based VLAs [44, 24, 6, 43]. The former uses discrete action tokens within the vision-language model’s vocabulary, enabling action generation similar to language modeling through next-token prediction. The latter leverages a policy head [5, 4, 26], such as a diffusion policy [44] or flow matching [6], to output continuous robot actions.
In this work, we employ the recently released DiffusionVLA [43] model, which integrates the Qwen2-VL [41] vision-language model with a diffusion model head for action prediction. Specifically, the model outputs language through the VLM backbone and the language embeddings are subsequently injected into the policy model to enhance policy learning. This operation allows the model to reuse reasoning information while achieving faster inference than ECoT [48]. In the following sections, we will discuss how we improve this VLA by enabling it to reason through robot affordances before selecting an action.
4 Methodology
This section introduces our approach to explicitly leveraging affordance as a foundation for the robot models. In Section 4.1, we formally define the concept of chain of affordance as a structured sequence of affordances, each representing an actionable insight. We provide detailed descriptions of four distinct types of affordances that together constitute this chain, explaining how each type contributes to understanding and executing complex tasks. In Section 4.2, we present two formats for representing the chain of affordances: a text format and an image format. We then discuss how these representations can be integrated into the policy learning process. Finally, in Section 4.3, we outline the pipeline used to automatically generate large-scale chain-of-affordance data.
Figure 1: An example of the chain of affordances in the Pour Tea task is shown here. The first row presents the text-based chain-of-affordance prompts, while the second row shows the image-based chain-of-affordance prompts.
4.1 Definition of Chain-of-Affordance
Given a dataset $\mathcal{D}=\{(\tau_{1},g_{1}),\dots,(\tau_{N},g_{N})\}$ consisting of $N$ expert demonstrations, where each demonstration $\tau_{i}$ is paired with a task description $g_{i}$ in natural language. Each task description $g\in\mathcal{G}$ specifies a composition of multiple sub-tasks, and each demonstration $\tau_{i}$ is represented by a sequence of observations. We define $z\in\mathcal{Z}$ as the affordance-based reasoning in natural language that guides the task. The model decomposes $z$ into four components, $z=\{z_{obj},z_{grasp},z_{spat},z_{move}\}$, where $z_{obj},z_{grasp},z_{spat},$ and $z_{move}$ represent object, grasp, spatial, and movement affordances, respectively. Our objective is to learn an intermediate language output $z:\mathcal{O}\times\mathcal{G}\rightarrow\mathcal{Z}$ that maps observations and task descriptions to affordance reasoning in natural language. This intermediate output provides specific guidance for action generation, enabling the generation of low-level actions $a\in\mathcal{A}$. Note that low-level actions are generated conditioned on the demonstration, task description, and affordance reasoning: $a\sim p(a|\tau,g,z)$.
Details for diverse robot affordance. In our approach, we model the robot affordance in the format of natural language as intermediate outputs. We will give detailed description on each type of affordance.
Object affordance: Object affordance enables the robot to determine what object to interact with and where it is located, particularly when a user query lacks explicit instructions. This foundational knowledge allows the robot to discern the object it should manipulate and localize it within its environment. In our setup, we require the robot to identify the object by name in natural language and predict its location in an image using a bounding box format. This identification provides the base knowledge for the robot’s decision-making and interaction processes.
Grasp affordance: Grasp affordance encompasses the possible functions or ways an object can be manipulated. This affordance goes beyond visual characteristics, linking observations directly to actions, and is crucial for tasks requiring 6-DoF (degrees of freedom) grasping [17, 40, 31]. Prior work has demonstrated the effectiveness of affordance prediction for stable object handling and placement. Representations of grasp affordance vary, including part segmentation or keypoints. In our work, we use a set of 2D points to represent the grasping point for an object.
Spatial affordance: Spatial affordance is about spatial relations as an indicator of a model’s ability to understand a 3D world. This affordance supports tasks like identifying free spaces for object placement or navigation. For instance, RoboPoint [47] locates free space, and SpatialVLM [8] predicts spatial relations quantitatively and qualitatively. In our work, we define an actionable destination for the robot, such as vacant space on a plate for placing objects or a clear path for cleaning. We represent spatial affordance as a set of 2D coordinates to indicate feasible areas for action.
Movement affordance: Movement affordance defines the trajectory a robot can follow during a task. This path may change depending on environmental factors, such as obstacles introduced along an intended trajectory. By modeling movement affordance, we provide the robot with adaptable paths for action, allowing it to respond dynamically to environmental changes and complete its task effectively. These affordances collectively enable the robot to understand and act upon various elements within its operational space, enhancing its interaction capabilities and responsiveness.
Updating affordance chain. During both the training and inference stages, only a subset of affordance is involved based on the robot and scene state. Specifically, the model first predicts object affordance and grasp affordance, which tells the model what and how to manipulate. After the object is captured by the robot, it predicts the spatial affordance. The movement affordance is predicted throughout the whole period of action. By updating the affordance chain in real-time, we ensure minimum computational overhead is introduced.
Notably, the language generation is optimized through next token prediction, and the action generation is updated via diffusion model objective. This technique is similar to visual grounding works. We utilized the pre-trained checkpoint of DiffusionVLA to initialize our model. All constructed data are co-trained.
4.2 Formatting of Chain-of-Affordance
In this section, we introduce two formats for representing the chain of affordance: one in text format and another in image format. The text-based format provides a structured, linguistic approach to affordance representation, while the image-based format offers a visual perspective that can enhance the interpretability of affordance cues in visually complex environments. Additionally, we describe a new module specifically designed to integrate the image-based chain of affordance prompts into the policy model, thereby enabling the model to leverage both visual and linguistic cues for action generation.
Text-based chain-of-affordance prompting. Natural language is the most common way to represent visual affordances. For example, in object affordance, a set of coordinates can specify the location of objects, while spatial affordance can be conveyed through another set of coordinates. Figure 1 (top) provides an example of using natural language to denote various affordances. In practice, we avoid using a fixed language template to describe affordances. Instead, we use ChatGPT to enrich language expression by varying formats, similar to data augmentation in natural language processing. This approach helps the model retain its conversational capabilities. Notably, our base model is inherently capable of reasoning in natural language, so we adhere to its architecture to process affordances in text format to the policy model.
Image-based chain-of-affordance prompting. Inspired by methods that enhance model performance through visual prompts, we introduce visual affordance augmentation, which supplements natural language descriptions with direct visual cues. This approach enables the model to interpret and utilize affordance information more effectively.
Specifically, we overlay coordinates or point trajectories directly onto the robot’s previous observation frame, serving as a form of chain-of-affordance prompting that visually guides the model. This overlay provides a clear, interpretable structure, linking objects and their potential actions within the scene, and helps bridge the gap between descriptive language and practical visual information. Examples of this approach are shown in Figure 1 (bottom), where affordances are visually layered onto the image.
For movement affordances, which are applicable across all scenarios, we use a thinner point trajectory. This design choice prevents these trajectories from drawing excessive attention. Additionally, we employ thicker and lighter colors to denote key elements such as the object’s bounding box, grasping points, and spatial affordances. These visual distinctions ensure that each affordance type is easy to identify and interpret at a glance, enabling the model to parse visual information effectively.
Injecting image-based chain-of-affordance prompting. Unlike text-based chain-of-affordance prompting, the base model does not natively support images as intermediate inputs to the policy network. To overcome this limitation, we propose an image chain-of-affordance injection module that incorporates visual knowledge directly into the policy model. Our approach involves augmenting the robot’s current observation with visual affordances, allowing the model to retain and utilize affordance-based cues directly from images. Specifically, we use the pre-trained visual encoder from base VLA to extract detailed feature representations from the image, breaking it down into multiple patches to capture a finer level of detail and spatial context. These visual tokens are concatenated and passed through Multi-Layer Perceptron (MLP) layers to reduce the channel dimensions. To inject the image chain-of-affordance prompting seamlessly, we leverage Feature-wise Linear Modulation (FiLM) [36], which enables adaptive conditioning of the model based on visual context. This branch is similar to the reasoning injection module in the DiffusionVLA [43]. By aligning image affordances with FiLM, we ensure that the model can interpret and respond to visual cues in its decision-making process effectively. The image-based affordance injection allows the model to process richer contextual information.
Figure 2: Robot setup and examples for real-world manipulation tasks. We evaluate seven real-world tasks on Franka robot arm equipped with two external Zed cameras and a Realsense 435i wrist camera.
4.3 Generating Chain-of-Affordance Data
To prevent overfitting on affordance diversity, generating large-scale, high-quality data is crucial. While the standard approach typically relies on direct human labeling, this method is both costly and labor-intensive. Therefore, in this section, we outline a comprehensive pipeline for automatically generating diverse affordances using a range of tools to streamline the data generation process.
Our pipeline begins with GPT-4o [35], which generates a detailed description of the scene and identifies relevant entities from the language instructions. This allows us to create contextually rich, entity-specific affordances tailored to the task at hand. Using these entities, we leverage Grounding DINOv2 [29] and SAM [22, 37] to produce bounding boxes around each identified object within the scene. SAM initially provides masks, which we convert into bounding boxes, and we finalize these by taking out the intersection over union (IoU) between the outputs of Grounding DINOv2 and SAM. This IoU-based refinement ensures that bounding boxes are accurately aligned with object contours. At this stage, we also capture the gripper’s position for subsequent affordance calculations.
To represent spatial affordance, we integrate RoboPoint [47], a state-of-the-art model that predicts spatial affordances directly within the image. Additionally, we prompt GPT-4o to annotate spatial points based on the scene context. The spatial predictions from RoboPoint and GPT-4o are then combined, after which we cluster these points to form a coherent representation, eliminating any outliers to maintain accuracy. This process captures spatial affordances that are not only visually aligned but contextually relevant.
For capturing movement trajectories, we employ CoTracker [18, 19], an advanced transformer-based tracking model. CoTracker enables us to follow the robot gripper’s path, recording its movement through the scene and gathering essential trajectory data. This movement data provides insight into how the robot interacts dynamically with the environment, adding temporal dimensions to our affordance representation. The combination of spatial affordances, grasping points, and movement trajectories enables us to model a rich, multi-faceted affordance landscape for each scenario. This automated, tool-assisted pipeline produces a detailed and diverse set of affordance annotations, significantly reducing the need for manual labeling.
5 Experiments
The objective of our experimental evaluations is to assess the efficacy of CoA as a reasoning framework that augments the policy-learning capabilities of a baseline robot foundation model. Specifically, we seek to address the following questions:
  • Does the robot foundation model demonstrate improved performance with the integration of CoA in real-world experiments?
  • What is the performance of CoA in simulation benchmarks?
  • How effectively does CoA generalize to challenging scenarios?
5.1 Evaluation on Real Robot
Experimental setup. We conduct evaluations using the 6-DoF Franka robot arm, a widely adopted setup for assessing generalizable robot policies. Our setup includes two third-person cameras (ZED cameras) positioned on either side of the robot arm, as well as an egocentric camera (Realsense D435i) mounted on the wrist. We design seven challenging tasks, such as pouring tea and cleaning the table, incorporating both short-horizon and long-horizon tasks to test the adaptability of our approach. We illustrate our robot setup and tasks with examples in Figure 2. Detailed descriptions of each task and the experimental setup are provided in the Appendix.
Train data. We utilize the Droid dataset [20] as an external data source. Specifically, since not all samples in Droid contain annotations for language instructions, we filter out these samples, resulting in a total of 39K trajectories. We then apply our pipeline to generate synthetic chain-of-affordance data for Droid and use this data for pre-training. Following this, we post-train the model on task-specific data, collecting 692 trajectories for seven tasks in total.
Table 1: Experimental Results for Multi-Task Learning. Our method achieved the best performance in both the in-distribution test setup and under visual changes..
Seven Tasks on Franka Robot Arm
Model $\setminus$ Tasks In-Distribution
CleanTrash PourTea NailHammer PlaceBread PlaceCar WipeWater HangCup Avg.
Diffusion Policy [9] 4/11 0 8/11 3/11 9/11 2/11 7/11 33/77 (42.93%)
Octo [33] 4/11 0/11 7/11 4/11 8/11 3/11 8/11 34/77 (44.13%)
OpenVLA [21] 4/11 0/11 9/11 6/11 9/11 5/11 9/11 52/77 (54.89%)
DiffusionVLA [43] 10/11 6/11 10/11 7/11 10/11 6/11 10/11 59/77 (76.60%)
CoA 10/11 8/11 9/11 10/11 10/11 9/11 10/11 64/77 (85.54%)
Model $\setminus$ Tasks Visual Generlization
CleanTrash PourTea NailHammer PlaceBread PlaceCar WipeWater HangCup Avg.
Diffusion Policy [9] 0/9 0/9 0/9 0/9 3/9 0/9 0/9 3/63 (4.76%)
Octo [33] 0/9 0/9 3/9 0/9 4/9 2/9 3/9 12/63 (19.05%)
OpenVLA [21] 3/9 0/9 0/9 4/9 4/9 3/9 0/9 14/63 (22.22%)
DiffusionVLA [43] 2/9 2/9 5/9 5/9 5/9 3/9 6/9 27/63 (42.85%)
CoA 4/9 4/9 6/9 6/9 5/9 4/9 7/9 36/63 (57.14%)
Table 2: Experimental Results for LIBERO benchmark. We report the success rate and standard error for four task suites.
LIBERO-Spatial LIBERO-Object LIBERO-Goal LIBERO-Long Average
Method / Task Success Rate ($\uparrow$) Success Rate ($\uparrow$) Success Rate ($\uparrow$) Success Rate ($\uparrow$) Success Rate ($\uparrow$)
Diffusion Policy [9] $78.3\pm 1.1\%$ $92.5\pm 0.7\%$ $68.3\pm 1.2\%$ $50.5\pm 1.3\%$ $72.4\pm 0.7\%$
ScaleDP [51] $79.1\pm 0.7\%$ $90.4\pm 0.9\%$ $73.6\pm 0.8\%$ $48.4\pm 1.2\%$ $72.9\pm 0.5\%$
Octo [33] $78.9\pm 1.0\%$ $85.7\pm 0.9\%$ $84.6\pm 0.9\%$ $51.1\pm 1.3\%$ $75.1\pm 0.6\%$
OpenVLA [21] $84.7\pm 0.9\%$ $88.4\pm 0.8\%$ $79.2\pm 1.0\%$ $53.7\pm 1.3\%$ $76.5\pm 0.6\%$
CoA $85.3\pm 0.9\%$ $93.1\pm 1.0\%$ $85.8\pm 0.9\%$ $55.0\pm 1.2\%$ $79.8\pm 0.5\%$
Baseline. We compare CoA with Diffusion Policy [9], Octo [33], OpenVLA [21], TinyVLA [44], DiffusionVLA [43]. The latter three methods are all vision-language-action models that achieve state-of-the-art performance in the real world. The DiffusionVLA is the same model our approach is built upon, but trained with vanilla reasoning instead of our proposed chain-of-affordance. To ensure a fair comparison, all models are fine-tuned on the same dataset we use for training our approach. All models are trained with the same number of iterations, and the last checkpoint is used for evaluation. It ensures that no model is cherry-picked for comparison.
Real robot experimental results. The experimental results are presented in Table 1. We assess the performance using two different settings: the in-distribution setup and visual generalization testing. In the in-distribution setup, our method surpasses all SOTA robot foundation models in terms of average success rate. Notably, our method exceeds the performance of OpenVLA by 30.65%, despite having a significantly smaller model size and less pre-training data. Compared to our baseline model, which employs vanilla reasoning, our method achieves a 14.29% increase in accuracy. To assess visual generalization, we further evaluate these advanced models under varying visual conditions. As observations become more complex—such as through the addition of distractors or vibrant backgrounds—the performance gap between CoA and OpenVLA/DiffusionVLA widens. These findings highlight the critical role of embodied, task-specific chain-of-affordance reasoning in enhancing the performance of VLA models.
5.2 Evaluation on Simulation
In this section, we examine the performance of CoA on LIBERO [27]. LIBERO is a robot learning benchmark comprising over 130 language-conditioned manipulation tasks. We follow the setting as in OpenVLA [21] open-sourced code and test on four task suites: LIBERO-Spatial, LIBERO-Goal, LIBERO-Object, and LIBERO-Long. The detailed experimental setup are in the Appendix.
Simulation experimental results. We compare with the Diffusion Policy [9], ScaleDP [51], Octo [33], and OpenVLA [21]. The experimental results of LIBERO are presented in Table 2. Our findings indicate that CoA consistently achieves superior performance across all evaluated settings, securing the highest success rate among the methods tested. Specifically, CoA achieves an overall success rate of 79.8%, outperforming OpenVLA, the previous best-performing method, by a margin of 3.3%. This improvement demonstrates the effectiveness of our approach in simulation. Compared to pre-trained robotic models like Octo and OpenVLA, CoA demonstrates non-trivial gains, showing its effectiveness in leveraging prior knowledge while adapting to diverse task requirements. Additionally, when benchmarked against train-from-scratch methods such as Diffusion Policy and ScaleDP, CoA exhibits substantial improvements, underscoring the advantages of our approach.
Figure 3: Spatial Affordance for CoA. CoA can identify free space for object placement..
Figure 4: Movement Generalization for CoA. CoA can avoid obstacles and operate safely.
Figure 5: Generalization on Object Pose. CoA can pick up objects with unseen poses, benefiting from grasp affordance.
5.3 More Experiments
Spatial affordance enables safe placement. We evaluate the effectiveness of our spatial affordance approach in the PlaceBread task, as illustrated in Figure 3. In this task, the robot is presented with a plate on which three distinct objects are already placed, and it is instructed to add a piece of bread onto the plate. Our method successfully identifies open areas on the plate, allowing it to accurately position the bread without interference, thereby enabling CoA to complete all three task scenarios with precision. In contrast, alternative methods, such as OpenVLA and DiffusionVLA, succeed in only one scenario each, failing to generalize across all spatial configurations. This experiment underscores the crucial role of spatial affordance in enhancing the model’s ability to recognize and utilize available space, ultimately improving task completion accuracy and reliability across varied setups.
Obstacle avoidance. Collision avoidance is essential for safe and effective physical interactions, as improper maneuvers can lead to significant damage or even catastrophic outcomes. We evaluated CoA’s obstacle avoidance capabilities in two specific tasks, illustrated in Figure 4. In the first task, we positioned a vase near the center of the task area to observe whether CoA could successfully navigate around the vase to retrieve a piece of paper trash. In the second task, we introduced a series of obstacles on a table, rearranging them in different configurations to assess the robot’s adaptability in maneuvering through varied layouts to complete the task. Our approach successfully completed all three scenarios, demonstrating robust collision avoidance and spatial adaptability. In contrast, OpenVLA failed to complete any of the tasks, while DiffusionVLA succeeded in only one scenario. These results highlight the critical role of integrating movement affordance into the model’s reasoning process, enhancing its ability to navigate complex environments and complete tasks with precision and safety.
Generalization to unseen object pose. We assessed CoA’s ability to generalize to previously unseen object orientations, as illustrated in Figure 5. Our evaluation focused on two objects: a hammer and a teapot. In the training phase, both objects were consistently presented with their handles oriented vertically relative to the robot. To test the model’s generalization capabilities, we introduced novel poses that were absent from the training data, challenging CoA to grasp these objects in unfamiliar orientations. We observed that CoA successfully managed most scenarios, demonstrating a remarkable ability to adapt to new object poses even without explicit training on these orientations. In contrast, OpenVLA succeeded only in the simplest cases, struggling with more complex orientations. However, when the objects were positioned horizontally relative to the robot, all models, including CoA, were unsuccessful in achieving a stable grasp. Despite this limitation, our grasp affordance approach shows promising results, enabling CoA to handle a wide range of novel object poses.
6 Conclusion
Explicit reasoning is essential for language models to handle complex tasks. In this work, we design a reasoning-aware foundation model for robotics, focusing on various affordances: object, grasp, spatial, and movement. These affordances form an interdependent chain: the robot identifies the target object and location, determines how to grasp it, decides where to place it, and navigates accordingly. By structuring this chain of affordances as intermediate language and image outputs and feeding this reasoning into the policy model, our Chain-of-Affordances (CoA) model outperforms baselines on real-world robotic tasks. CoA models also generalize well to complex environments, tackling challenges like grasping in unfamiliar orientations, avoiding obstacles, and spatial generalization. Our approach provides a novel perspective on designing reasoning chains to enhance embodied control in real-world settings.
Supplementary Material
7 Video Demo
We provide a video recording in the supplementary material.
Figure 6: Visual Generalization. We evaluate each method on multi-task learning and visual generalization, which includes handling additional distractors and interference from colored light. We also test the ability to grasp objects of the same type but with varying shapes, such as different teapots, as well as teapots in different orientations.
Figure 7: More detailed examples of successful Chain-of-Affordance.
Table 3: Summarization for the number of demonstrations and average trajectory length for our real-world tasks.
# Task # of Demonstrations Average Trajectory Length
1 PlaceCar 89 301.8
2 PlaceBread 102 113.2
3 NailHammer 80 182.8
4 PourTea 91 429.4
5 CleanTrash 185 114.3
6 WipeWater 62 199.9
7 HangCup 83 131.4
8 Evaluation Tasks
In this section, we give a detailed description of the evaluated tasks that we discussed in Table 1. We provide the number of demonstrations for each task and the average trajectory length in Table  3.
  • PlaceCar. We randomly place the toy car on the right side of the drawer. The model is asked to pick up the toy car, put it into the drawer, and eventually close the drawer. This is a long-horizon task that requires multiple steps of action.
  • PlaceBread. The model needs to pick up the bread and place it on an empty spot on the plate, avoiding placing it on the fruit. The bread is randomly placed on the table. The model needs to pick up the bread and place it on the empty spot on the plate.
  • NailHammer. We evaluate the model’s proficiency in utilizing tools effectively by assessing its ability to perform a sequence of precise actions with a hammer. The model must first identify the correct grasp point on the hammer, ensuring a stable and ergonomic grip suitable for controlled operation. It must then carefully pick up the hammer without causing it to topple or disturb its surroundings. Once the hammer is securely held, the model is tasked with driving a nail into a designated spot with precision.
  • PourTea. In this task, the robot is required to perform a sequence of actions involving a tea cup and a teapot. First, the robot must place the tea cup onto the tea tray. Next, it needs to pick up the teapot and pour tea into the teacup. Both the tea cup and the teapot are randomly positioned within a defined range on the table. A key aspect of the task is the robot’s ability to accurately grasp the teapot by its stem. To ensure consistency during data collection, the tea pot’s stem is always oriented facing the robot, simplifying the grasping process while still challenging the model’s precision and manipulation skills.
  • CleanTrash. In this task, the robot is required to perform a sequence of actions to clean up trash on a table. The task has two distinct scenarios. In the first scenario, with no obstacles, the robot must identify and pick up the randomly placed trash, then deposit it into the trash bin. The trash items are distributed across the table in a random manner. In the second scenario, a flower pot is placed on the table as an obstacle. The robot must avoid colliding with the flower pot while picking up the trash and placing it into the trash bin. The trash’s location remains random, and the robot must navigate carefully to avoid knocking over the flower pot during the cleanup process. A key aspect of this task is the robot’s ability to accurately avoid the flower pot while maintaining efficiency in picking up and discarding the trash.
  • WiperWater In this task, the robot is required to clean up water from a table by using a sponge. The sponge is placed on the right side of the table, and the robot must pick it up and use it to wipe the water from the surface, moving from right to left. During this process, the robot must avoid any objects placed on the table, such as vases, cups, boxes, and other items. A key challenge in this task is the robot’s ability to manipulate the sponge effectively while navigating around the obstacles without causing any collisions, ensuring that the entire table is cleaned efficiently. The robot’s precision in both grasping the sponge and avoiding the table items is critical for completing the task successfully.
  • HangCup In this task, the robot is required to pick up cups that are randomly scattered on the table and hang them on a cup rack. The robot must handle the cups carefully to avoid damaging them and ensure that the rack is not disturbed or knocked over during the process. The task challenges the robot’s precision in both grasping the cups and placing them securely on the rack while maintaining stability in the environment. Successful completion relies on careful manipulation and accurate placement.
Setup for visual generalization. In this scenario, we evaluate the model’s robustness and its ability to generalize visual perception across diverse and challenging environmental conditions. The robot is tasked with performing manipulation tasks while navigating visual complexities such as randomly placed distractors, varying lighting conditions, and a visually cluttered, colorful background. These challenges are designed to test the model’s capability to stay focused on the primary task, effectively filter out irrelevant visual distractions, and adapt to dynamic and unpredictable visual environments. The objective is to ensure the robot can consistently and accurately identify and interact with target objects, even under significant deviations from typical operational settings.
9 Details for Real Robot Experiments
We train our method in a multi-task setting without relying on pre-trained weights from DiffusionVLA. Instead, we leverage our constructed dataset for pre-training. Specifically, we initialize the learning rate at 2e-5 and maintain a fixed learning rate throughout the pre-training phase, which spans 5 epochs. During this stage, the parameters of the pre-trained Vision-Language Model (VLM) are frozen, and LoRA is employed to fine-tune the model. For fine-tuning, we adopt a similar approach, starting with an initial learning rate of 2e-6. However, in this phase, we apply a cosine learning rate decay schedule and train the model for an additional 5 epochs. This training strategy ensures both effective adaptation and stability across pre-training and fine-tuning stages, optimizing the model for multi-task performance.
For the baselines, we generally adopt a consistent training strategy. In the case of OpenVLA, the vanilla implementation utilizes only a single camera view. To extend this, we incorporate all three camera views, feeding each view into the same visual encoder and concatenating their outputs for processing. We leverage OpenVLA’s pre-trained weights and trains for 20 epochs, as we observe that it typically requires a longer training time to achieve convergence. For the Diffusion Policy, we utilize DistilBERT to process language instructions, following an approach similar to YAY [38]. As for DiffusionVLA, we employ their pre-trained weights and construct a reasoning dataset using their data construction pipeline to maintain consistency with their methodology. To ensure fair evaluation, we use the final checkpoints of all models, including ours, avoiding any form of cherry-picking. This approach allows for a robust comparison and highlights the performance differences across the various models.
10 Details for LIBERO Simulation
LIBERO is a robot learning benchmark comprising over 130 language-conditioned manipulation tasks. We follow the setting as in OpenVLA [21] open-sourced code and test on four task suites: LIBERO-Spatial, LIBERO-Goal, LIBERO-Object, and LIBERO-Long.
Each suite includes 10 distinct tasks with 50 demonstrations per task. Each task suite emphasizes unique challenges in imitation learning: LIBERO-Goal features tasks with similar object categories but different goals. LIBERO-Spatial requires policies to adapt to varying spatial arrangements of the same objects. LIBERO-Object keeps the layout consistent while changing the objects. During experimentation, our method uses a static camera, and a wrist-mounted camera all methods are evaluated across 1500 trials in total. We filter out the failure data and increase the image resolution to $224\times 224$. The affordance data is generated using our proposed pipeline for data in LIBERO. In Table 2, we directly cite the results of Diffusion Policy, Octo, and OpenVLA from OpenVLA’s paper. Therefore, to ensure all methods are evaluated fairly, we evaluated our methods across 500 trials for each task suite, and the reported performance is the average success rate over three random seeds. We use the same test data as in OpenVLA. For the baseline ScaleDP, except for using all camera views, all other implementations kept the same.